OpenStudy (anonymous):

FAN & MEDAL TO ANYONE WHO HELPS Find the value of the following expression: (28 ⋅ 3−5 ⋅ 60)−2 ⋅ 3 to the power of negative 2 over 2 to the power of 3, whole to the power of 4 ⋅ 228 (5 points) Write your answer in simplified form. Show all of your steps. (5 points)

OpenStudy (anonymous):

i have no clue how to do this

OpenStudy (anonymous):

did this get deleted?

OpenStudy (anonymous):

yeah bc it was in the english section.i deleted it @golfing

OpenStudy (anonymous):

i have no clue how to read it use the equation editor below to make it legible

OpenStudy (anonymous):

okay, it might take a second XD

OpenStudy (anonymous):

\[(2^{8} \times 3^{-5} \times 6^{0}) ^{-2} ( \frac{ 3^2 }{ 2^3 } \times 2^{28}\]

OpenStudy (anonymous):

on the fraction, the 3^2 is actually 3^-2 ( negative 2)

OpenStudy (anonymous):

\[6^0=1\] so ignore that part

OpenStudy (anonymous):

\[(2^8\times 3^{-5})^{-2}\] multiply the exponents at get \[2^{-16}3^5\]

OpenStudy (anonymous):

\[\frac{3^{-2}}{2^3}=\frac{1}{3^22^3}\]

OpenStudy (anonymous):

how are you doing this

OpenStudy (anonymous):

\[b^{-n}=\frac{1}{b^n}\]

OpenStudy (anonymous):

and \[(b^n)^m=b^{m\times n}\]

OpenStudy (anonymous):

how do i write that in words

OpenStudy (anonymous):

that is entirely up to you

OpenStudy (anonymous):

idk how to do that